OpenStudy (loser66):

How to get it? \[\large \sum\limits_{k\geq n} \dfrac{1}{2^k}= \dfrac{1}{2^{n-1}}\] Please help

OpenStudy (gorv):

can type it up again plzzz

OpenStudy (gorv):

its not readable

OpenStudy (loser66):



OpenStudy (loser66):

I read but don't know how to get this part.

ganeshie8 (ganeshie8):

isnt it just a geometric series ?

OpenStudy (loser66):

yes, but we have formula to get it as = \(a+ar+ar^2+.....+ar^n =a\dfrac{1-r^{n+1}}{1-r}\) for 0 <r <1

ganeshie8 (ganeshie8):

\[\large S_{\infty} = \dfrac{a}{1-r}\]

ganeshie8 (ganeshie8):

infinite series formula requires |r| < 1 but your partial sum formula doesn't require it

OpenStudy (loser66):

the last term is ar ^(n-1)

ganeshie8 (ganeshie8):

there is no last term, its an infinite series k >= n means, k = m to infinity

OpenStudy (loser66):

yes, I understand it

OpenStudy (gorv):

but ganesh i have little doubt dont mind

OpenStudy (gorv):

here our first term will be 2^n right not 1???

ganeshie8 (ganeshie8):

yes a = 1/2^n r = 1/2 right ?

OpenStudy (loser66):

yes

OpenStudy (gorv):

yeah i mean that...okkkk fine :)

OpenStudy (gorv):

just conforming ..that we r goin to start with n :P

OpenStudy (loser66):

But in this stuff, it doesn't last long to infinitive, it starts from m and stop at m+1 then take sum of those part give us ... something like above, but I don't get how to get 1/2^(n-1)

OpenStudy (gorv):

n will be a real number

OpenStudy (gorv):

but upper limit to k is not given

ganeshie8 (ganeshie8):

any partial sum is less than the infinite sum because the series is increasing

OpenStudy (loser66):

Got you!! make perfect sense to me to skip the part I don't get.

ganeshie8 (ganeshie8):

\[\large \sum\limits_{k\ge m} \dfrac{1}{2^k} ~\gt ~~ \sum\limits_{k= m}^{m+k-1} \dfrac{1}{2^k} \]

OpenStudy (loser66):

next?

ganeshie8 (ganeshie8):

they are using that

ganeshie8 (ganeshie8):

In the attached doc : the partial sum in line3 is less than tha infinite sum in line4

ganeshie8 (ganeshie8):

its like saying this : 10^2 = 10^2 < 10^3 < 10^4 = 10000

OpenStudy (loser66):

I have class now, I have to go. Thanks for the help

ganeshie8 (ganeshie8):

the end result is that 10^2 < 10000

ganeshie8 (ganeshie8):

Okay, its simple for you if u look at it calmly :)

OpenStudy (loser66):

Again, thanks a lot